OpenStudy (anonymous):

Find the sum of the first 9 terms of the sequence. 2, -7, -16, -25, ...

hartnn (hartnn):

do you see any pattern in your sequence ?

OpenStudy (anonymous):

Well the difference between the three numbers to the right are all 9, but the difference between 2 and -7 is only 5 sooo I'm a little confused. @hartnn

OpenStudy (anonymous):

How do I figure out what the other numbers are?

hartnn (hartnn):

the difference between 2 and -7 : \(\Large 2 - (-7) = 2+7 = 9 \\ \large \ddot \smile \)

OpenStudy (anonymous):

Ohhhhhh! So it would be: \[2, -7, -16, -25, -34, -43, -52, -61, -70 = -306\]

OpenStudy (anonymous):

I feel like that's wrong though

OpenStudy (anonymous):

Wouldn't I need to make some of the numbers positive? Would numbers -34 and -70 actually be positive? @hartnn

hartnn (hartnn):

why you need to make them positive ?? since we have a common difference of -9 , this sequence is arithmetic sequence. Have you taught how to find sum of terms in arithmetic sequence ? any formula ?

OpenStudy (anonymous):

Formula? No. I'm just supposed to add them all together once I know what the numbers are, right? So is my answer of -306 wrong?

hartnn (hartnn):

There is an easier way to do it, using formula : \(\large S_n = (n/2)[2a+ (n-1)d]\) where, n= number of terms a = 1st term d = common difference (imagine if you're asked sum of 1000 terms...would you literally add 1000 terms ?)

hartnn (hartnn):

for your problem, first term = a = 2 d= common difference = -9 and since we need sum of 9 terms , n = 9

hartnn (hartnn):

And if you apply that formula, you get -306 indication you're anyways corrrect! :)

OpenStudy (anonymous):

Cool! Thank you so much! @hartnn

hartnn (hartnn):

welcome ^_^